OpenStudy (commonlizer):

Two equations are given below: m + 3n = 10 m = n − 2 What is the solution to the set of equations in the form (m, n)? A (1, 3) B (2, 4) C (0, 2) D (4, 6)

OpenStudy (commonlizer):

my best guest is C

rebeccaxhawaii (rebeccaxhawaii):

okay so m=n-2 plug it into the first equation

rebeccaxhawaii (rebeccaxhawaii):

(n-2)+3n=10

rebeccaxhawaii (rebeccaxhawaii):

n-2+3n=10 n+3n=10+2 get n's on one side whats n+3n=? and 10+2=?

OpenStudy (commonlizer):

10 + 2 = 12 and I think that 3n is cancelled right

rebeccaxhawaii (rebeccaxhawaii):

no 3n+1n=?

OpenStudy (commonlizer):

oh okay so 3+1 = 4

rebeccaxhawaii (rebeccaxhawaii):

3n+1n=4n

OpenStudy (commonlizer):

so I am taking it is D right

rebeccaxhawaii (rebeccaxhawaii):

okay so 4n=12 divide both by 4 so n=3

rebeccaxhawaii (rebeccaxhawaii):

please stop guessing, im taking you through the problem

OpenStudy (commonlizer):

I am sorry plz continue

rebeccaxhawaii (rebeccaxhawaii):

okay now to get m, we plug in 3 for n in one of the equations m=n-2 so plug in n=3 m=3-2

rebeccaxhawaii (rebeccaxhawaii):

whats 3-2?

OpenStudy (commonlizer):

1

OpenStudy (chloegauv):

The answer C

rebeccaxhawaii (rebeccaxhawaii):

correct so m=1 and n=3 so they put in it (m,n) order (1, 3)

OpenStudy (chloegauv):

m=0, n=2

OpenStudy (commonlizer):

thank you all even you @rebeccaxhawaii for taking me through and help me out even when I was guessing

rebeccaxhawaii (rebeccaxhawaii):

no problem c: good luck with your studies.

OpenStudy (commonlizer):

you to lol